OpenStudy (anonymous):

5/8 x+1/16 x=3/16+x Accidentally deleted this one...

OpenStudy (saifoo.khan):

brackets! :/

OpenStudy (anonymous):

There weren't any on it when i copied it

OpenStudy (anonymous):

multiply by the lcd 16[(5/8)x+(1/16)x=3/16 + x] 10x+x=3+16x -4x=3 x=-3/4

OpenStudy (anonymous):

x=-0.6

OpenStudy (anonymous):

1) Multiply by 16 (16) * (5/8) x + (16) * (1/16) x = (16) * (3/16) +(16)*x = 10x + x = 3 + 16x 2) Group like terms -3 = 5x x = -3/5

OpenStudy (anonymous):

yes, -3/5 , sorry